OpenStudy (anonymous):

What is the solution of the system? Use substitution. y = 6x + 2 3y - 18x = 12

OpenStudy (blake57roger):

a

OpenStudy (anonymous):

there is no solution the two graphs are parallel

OpenStudy (a1234):

y = 6x + 2 3y - 18x = 12 3(6x + 2) - 18x = 12 18x + 6 - 18x = 12 6 = 12? No solution.